A cellular phone company offers a contract for which the cost C, in dollars, of t minutes of telephoning is given by C=0.25(t-65.95), where it is assumed that t>=300 minutes. What times will keep cost between $108.95 and $131.45?
For the cost to be between $108.95 and $131.45, the telephoning time must be between ___ minutes and ___ minutes.

Answer by Theo(2978)

take your original equation of
c = .25 * (t - 65.95) and solve for t.
divide both sides by .25 to get
c/.25 = t - 65.95
add 65.95 to both sides of the equation to get
(c/.25) + 65.95 = t
just plug in the values for the low c and the high c and you get the time in minutes that would be minimum and maximum.
for the minimum, the equation becomes:
(108.95 / .25) + 65.95 = t
use parentheses to make sure the order of the arithmetic is clear.
simplify to get
435.8 + 65.95 = 5
simplify further to get
t = 501.75
that's your low number of minutes
solve the equation again using c = 131.45 to get the high end minutes.